OpenStudy (seattle12345):

How do I solve this?: A man has 22 coins in his pocket, all of which are dimes and quarters. If the total value of his change is 355 cents, how many dimes and how many quarters does he have?

OpenStudy (mertsj):

Let x = number of dimes 22-x = number of quarters 10x+25(22-x)=355

OpenStudy (lgbasallote):

let d = dime q = quarters d + q = 22 since dime = 10 cents and quarters = 25 cents 10d + 25q = 355

OpenStudy (mertsj):

Solve the equation.

OpenStudy (anonymous):

9 Quarters and 13 Dimes

OpenStudy (mertsj):

yep

OpenStudy (seattle12345):

10x+25(22-x)=355 10x + 550 – 25x = 355 -15x + 550 = 355 -15x = 195 X= -13

OpenStudy (mertsj):

@seattle12345 -15x=-195 x=13

OpenStudy (seattle12345):

I might be missing something but how did you get -195?

OpenStudy (seattle12345):

never mind I see, should be adding -550 to both sides lol. thanks for your help,,. I have trouble translating word problems into equations.

OpenStudy (seattle12345):

@lgbasallote, or anyone else, If you don't mind can you show me how I would use substitution?

OpenStudy (lgbasallote):

d + q = 22 10d + 25q = 355 by substitution: note: d = 22 - q 10d + 25q = 355 10(22-q) + 25 q = 355 220 - 10q + 25 q = 355 220 + 15q = 355 15q = 355 - 220 15q = 135 q = 9 note: d = 22 - q d = 22 - 9 d = 13

OpenStudy (seattle12345):

thanks guys you helped me out alot.